# StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land-Surface Predictions

### v2

@@ Appendix C Experimental Details
StefaLand with resConn 30 2 NVIDIA V100 80 GB
StefaLand without resConn 26 2 NVIDIA V100 80 GB
@@
StefaLand-no resConn HBV1.1 300 2 NVIDIA V100 80 GB
  • Note: All values except pretraining are for the CAMELS benchmark experiment. The relative differences in computational requirements are consistent across other experiments.
C.1 Pretraining Data Handling
To handle large volumes efficiently, data were stored in shards rather than fully loaded into memory. At the start of each epoch, shards were randomly reindexed, providing diverse samples while avoiding costly materialization of the full dataset. This strategy ensures both reproducibility and wide sample coverage. The accompanying code release in the reproducibility statement enables reconstruction of this dataset and reproduces the full pretraining pipeline.
C.2 Dataset Splitting
@@
Appendix D Model Architectures
Figure D1: StefaLand model architecture with residual connections. The pretrained StefaLand encoder (frozen during fine-tuning) processes static attributes and generates embeddings that are combined with meteorological forcings through residual connections. This architecture enables iterative integration of transformer features with input data through the CNN module, allowing the model to effectively leverage pretrained representations while adapting to task-specific requirements.
Figure D2: StefaLand model architecture without residual connections. In this configuration, the frozen transformer embeddings are used only once through a standard adapter that combines them with forcings and static features. This represents a more conventional fine-tuning approach where transformer features are not integrated iteratively throughout the decoding process.
Figure D3: Different adapter architectures tested in our experiments. Left: Gated Adapter with a selector mechanism that controls information flow. Center: BottleNeck Adapter with compression and expansion phases. Right: Residual Adapter that adds transformer features through a skip connection.

### v3

@@ C.3 Computational Resources
StefaLand ResConn 30 2 NVIDIA V100 80 GB
StefaLand no Adapter 26 2 NVIDIA V100 80 GB
@@
StefaLand-no Adapter HBV1.1 300 2 NVIDIA V100 80 GB
  • Note: All values except pretraining are for the CAMELS benchmark experiment. Relative computational differences are consistent across other experiments.
C.4 Pretraining Data Handling
Figure 4: Spatial distribution of the global streamflow dataset. Basins are categorized according to the availability of runoff observations: basins with relatively abundant runoff records (blue), basins with sparse runoff records (orange), and basins without runoff data (green). Marker size corresponds to basin area, classified into three categories based on the 33rd and 67th percentiles of catchment areas.
A global dataset was constructed for model pretraining, including 8,634 catchments and designed to characterize climatic, ecological, soil, topographic, geological, and socioeconomic conditions. The dataset includes both daily meteorological forcings and long-term averaged static attributes. Daily meteorological variables comprise precipitation, downward shortwave radiation, relative humidity, maximum temperature, and minimum temperature, derived from the Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP) datasets at a spatial resolution of 0.1° (Beck et al., 2022, 2019). Potential evapotranspiration was estimated using the Hargreaves method.
The study catchments are divided into three groups: 3,434 GRDC catchments with relatively abundant historical runoff records, 3,248 GRDC catchments with sparse runoff records, and 1,952 HydroBasins level-8 catchments without runoff observations (GRDC, 2024; Lehner and Grill, 2013). Ecosystem states are represented by forest and grassland cover fractions derived from the Climate Change Initiative (CCI) land cover dataset (ESA, 2017), along with the Normalized Difference Vegetation Index (NDVI) from MODIS (Didan, 2015a). Soil properties include sand, silt, and clay fractions from the Harmonized World Soil Database (HWSD) (FAO et al., 2012).
Topographic attributes include elevation, slope, and aspect obtained from the Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011; Ramcharan et al., 2018), as well as terrain-derived soil depth from the Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers dataset (Pelletier et al., 2016). Geological attributes comprise carbonate sedimentary rock fractions from the Global Lithological Map (GLiM) (Hartmann and Moosdorf, 2012) and rock porosity and permeability from the GLobal HYdrogeology MaPS (GLHYMPS) dataset (Gleeson et al., 2014).
Socioeconomic conditions are characterized using population density from the Gridded Population of the World (GPW) v4 dataset (CIESIN, 2016), gross domestic product and population data from the gridded global GDP and Human Development Index datasets (Kummu et al., 2018), and forest intactness from the Intact Forest Landscapes dataset (Potapov et al., 2017). All static attributes were mapped to a common 0.01° grid prior to basin-scale aggregation to ensure spatial consistency across data sources and improve spatial averaging over irregular basin geometries. A complete list of variables is provided in Table 19.
C.5 Dataset Splitting
@@
Appendix D Additional Figures
Figure 5: Fine-tuning pipeline used in our downstream experiments. Static attributes and meteorological forcings are encoded by the pretrained StefaLand encoder (frozen), then passed through a task adapter and sequence model (LSTM), followed by projection layers to generate predictions optimized with a task loss against ground truth.
Figure 6: Adapter architectures evaluated in our experiments. We compare a gated adapter, a bottleneck adapter with compression and expansion stages, A mixture of Experts and a basic feedforeward and a residual adapter that injects pretrained features via a skip connection.